# Visually Prompted Benchmarks Are Surprisingly Fragile

### v1

@@ Abstract
A key challenge in evaluating VLMs is testing models’ ability to analyze visual content independently from their textual priors. Recent benchmarks such as BLINK probe visual perception through visual prompting, where questions about visual content are paired with coordinates to which the question refers, with the coordinates explicitly marked in the image itself. While these benchmarks are an important part of VLM evaluation, we find that existing models are surprisingly fragile to seemingly irrelevant details of visual prompting: simply changing a visual marker from red to blue can completely change rankings among models on a leaderboard. By evaluating nine commonly-used open- and closed-source VLMs on two visually prompted tasks, we demonstrate how details in benchmark setup, including visual marker design and dataset size, have a significant influence on model performance and leaderboard rankings. These effects can even be exploited to lift weaker models above stronger ones; for instance, slightly increasing the size of the visual marker results in open-source InternVL3-8B ranking alongside or better than much larger proprietary models like Gemini 2.5 Pro. We further show that low-level inference choices that are often ignored in benchmarking, such as JPEG compression levels in API calls, can also cause model lineup changes. These details have substantially larger impacts on visually prompted benchmarks than on conventional semantic VLM evaluations. To mitigate this instability, we curate existing datasets to create VPBench, a larger visually prompted benchmark with 16 visual marker variants. VPBench and additional analysis tools are released at https://lisadunlap.github.io/vpbench/.
Figure 1: Small, seemingly irrelevant changes in visual prompting dramatically alter VLM predictions. Left: Qwen2.5-VL accuracy under different visual marker variants. Changes in marker size, shape, color, and label position lead to significant accuracy swings up to 13%. Right: such variations can reorder leaderboards, with model rankings shifting even when nothing about the underlying task changes.
@@ 1 Introduction
We explore three such sources of evaluation instability across eight modern VLMs on BLINK [9] as well as two more large scale datasets, VPBench-RD and VPBench-SC, that we curate using DA2k [26], and SPair [16]. First, we demonstrate the effect of sample choice: random resampling of image subsets, matched in its modest size to BLINK and drawn from a fixed pool of visually prompted tasks, leads to substantially reordered model rankings, despite the subsets being statistically indistinguishable in size and difficulty. Second, we examine visual prompt formatting. On all dataset , we evaluate our models with 16 different marker styles varying in size, shape, color, and label placement. We find that marker style can cause accuracy swings of up to 21% on the same image-question pairs, often causing ranking reversals among state-of-the-art models. Finally, we show that inference-time implementation details which are imperceptible to humans such as JPEG compression further perturb results in statistically significant ways. Additionally, we find that this is specific to visually prompted tasks, as applying the same intervention to more traditional VLM benchmarks does not significantly change the results. We also demonstrate how this fragility can be exploited to “game” leaderboards. For example, strategically selecting the visual marker to be a square instead of a circle causes a weaker model like InternVL3-8B to rank above stronger models like Gemini 2.5 Pro on the BLINK relative depth estimation task.
These findings suggest that much of the variation among model performance reported on vision-language benchmarks comes not from differentiated intrinsic capabilities of grounding language in vision, but from incidental details of prompting, sampling, and implementation. To address this, we release the expanded visually prompted datasets as VPBench. VPBench is a benchmark covering 16 different visual marker variants, and overall boosting the dataset size from 224 samples in BLINK relative depth and semantic correspondence to 35,088 annotated images across relative depth (VPBench-RD) and semantic correspondence (VPBench-SC). Lastly, we provide suggestions on how to more robustly evaluate visually prompted tasks and guidance on when to trust the leaderboard rankings. We release our proposed VPBench along with our inference code, which supports varying visual markers and image compression settings, as a reference inference for stable evaluation at https://lisadunlap.github.io/vpbench/.
@@ 3 Visually Prompted Tasks and Probing Them
Figure 2: Examples of visually prompted tasks. Visually prompted tasks (VPTs) involve placing visual markers in the image to ask questions such as relative depth and semantic correspondence.
@@ 3.1 What are Visually Prompted Tasks?
A visually prompted task explicitly marks regions of an image and asks about relationships within or between images [9]. We picked two typical visually prompted tasks (Fig. 2). In semantic correspondence, two images with marked points are shown and the model is asked which marker corresponds to the same object part or region as the reference marker. In relative depth, two locations in an image are marked with “A” and “B,” and the model is asked which is closer to the camera.
@@ 3.2 Experimental Setup
The BLINK dataset [9] systematizes these visually prompted tasks via curated image–question pairs with region-level markings. It contains relative depth and correspondence tasks with high annotation quality. However, we noticed that within BLINK, visual markers are not standardized: some examples use dots, others boxes or arrows. When these seemingly cosmetic differences are altered, we noticed non-trivial change in model performance. This motivated our investigation into how VLM benchmarks may be fragile given seemingly irrelevant design choices.
As will be detailed in Section 4, BLINK’s small size, approximately 100 examples per split, makes it difficult to separate true performance variation from sampling noise. To reduce this ambiguity and enable more robust analysis, we curate a new benchmark called VPBench. VPBench is sourced from two larger datasets that follow the same visually prompted task format: DA2K [26] and SPair-71k [16], with the former used for relative depth estimation (VPBench-RD) and the latter for semantic correspondence (VPBench-SC). DA-2K, originally introduced for depth annotation, contains thousands of images with dense geometric labels; we repurposed it into a relative-depth task by converting pixel-level depth into pairwise comparisons following BLINK’s protocol. SPair-71k (SPair), designed for semantic correspondence, provides tens of thousands of image pairs with detailed keypoint matches; by adopting BLINK’s prompting style, we frame SPair as a visually prompted correspondence benchmark that bridges traditional CV evaluation with modern VLM usage. We then apply all interventions for the following sections to all 4 datasets.
@@ 4 Statistically Equivalent Sampling Yields Different Results
Why sampling matters here?
This sensitivity arises because BLINK, like many newly-released multimodal benchmarks, deliberately targets unsaturated capabilities i.e. tasks where performance levels are far from ceiling, typically in the 30–60% range rather than the 80–90% common in relatively mature VLM tasks, consequently, these VPT accuracy variation would be much higher than knowledge-focused ones like MME (See Fig. 5). At such accuracy levels and sample size, even a few items can noticeably shift results: in BLINK, a 3% accuracy change corresponds to only two or three additional correct responses, which in a multiple-choice format could easily arise by chance. The resulting variance is reflected in the confidence intervals in Figure 3, and we further highlight its impact by showing how results fluctuate as the dataset size decreases.
@@ 5 Visual Marker Styles Shuffle Leaderboards
We investigate whether the style of the visual marker used in a prompt influences model performance, and whether it can change the relative ordering of models’ accuracy. This question is motivated by an observation of inconsistent marker styles in the BLINK benchmark dataset [9]. The visual prompts BLINK employs as part of the question context occasionally switched in color or text positioning and we noticed, for example, that simply switching a marker’s color from red to blue questions led to measurable changes in a model’s accuracy on BLINK, suggesting current models are over-reliant on specific visual cues. In this section we create a set of marker variants and study their effects on model performance. We report results on VPBench in the main paper and additionally report results for the BLINK subsets in the Appendix.
@@
Figure 6: Small marker changes cause large, model-specific accuracy shifts and rank shuffles. Rank (top) and accuracy (bottom) variations on (a) VPBench-RD relative depth and (b) VPBench-SC semantic correspondence across 16 visual marker variants (full chart in the Appendix). Rows are models and columns are marker styles; each cell reports the change in accuracy relative to the default BLINK marker, which uses a red circle with the label above. Positive values indicate improvements and negative values indicate degradations.
@@
Based on the most common marker seen in BLINK relative depth and semantic correspondence, we define a default marker style to be a small red circular marker with a numeric label placed above. We then create 16 alternative styles (commonly seen in visual prompting literature) spanning 4 categories: color (blue instead of red), shape (square instead of circle), size (larger marker, radius increased from the default to size 10), and text position (label moved to below the marker). Figures 6 and 6 display the change in accuracies and rankings of each model on relative depth and semantic correspondence tasks.
Figure 7: Impact by marker type. Mean absolute accuracy shift on VPBench-RD induced by each marker variant.
@@
To determine whether observed accuracy differences between markers are statistically significant rather than due to sampling variability, we draw $B$ bootstrap replicates $\mathcal{D}^{s}$ by sampling $N$ items with replacement. On each replicate, we compute accuracies with the default and an alternative marker and take the paired difference
@@
Figures 6 and Figure 6 show the model accuracy on the default visual marker (red circle, text at the top) along with the delta in performance seen across 16 different visual markers of varying sizes, shapes, colors, and positions. We see that if you change the visual marker, the takeaways about which model is superior can be vastly different. Beyond absolute accuracy, marker changes often altered the relative performance of models. For example, Llama 4 Scout outperformed Gemini 2.5 Flash under the default red marker for relative depth, but under the blue marker condition their ranks reversed, with Gemini 2.5 Flash scoring as the highest performing model. We also see cases of incredibly large differences in accuracy, such as Llama 4 Scout and InternVL for semantic correspondence, where making the text size of the marker smaller results in a 10+% drop in accuracy and dropping Llama from rank 3 to rank 6.
@@
Figure 7 displays the average magnitude of accuracy difference across each marker for the relative depth task, altering size or label position generally produced larger effects on accuracy than changing color or shape. Changing the marker’s color (red to blue) had a large impact on certain models, hinting that those models might be overfit to the color distribution of markers seen in their training or the benchmark (since BLINK’s default is red, some models may have learned to specifically attend to red circles as a prompt cue). Changing the marker’s size and shape or changing the text marker from A/B to 1/2 had a measurable but less prominent effect on accuracy. Notably, we did not find a single marker style that was universally best or worst for all models. These idiosyncratic responses point to each model having its own biases in visual prompt processing.
@@
We explicitly demonstrate that visually prompted leaderboards can be gamed by jointly selecting a marker style and an i.i.d. split that favor a particular model, see Figure 8. On BLINK relative depth, for example, we can lower internVL3 from rank 4 to rank 8 by simply changing the marker to be a square, and can similarly raise its rank to 3, higher than Gemini 2.5 Pro, by increasing the font size of the marker. This underscores that, without standardized marker conventions, visually prompted evaluations can give misleading impressions of model ability.
Figure 8: Manipulating leaderboards for InternVL3-8b. Performance comparison: optimizing for InternVL3-8B’s ranking on BLINK relative depth. Leaderboard rank can be shifted by benign visual prompt choices rather than changes in model capability.
@@ 6 Imperceptible differences matter as well
Figure 9: Change in rank seen across JPEG compression rates. Individual dots with no lines indicate models whose rank remained constant across all compression levels. JPEG compression quality substantially reorders model rankings on visually-dependent tasks (BLINK RD), whereas rankings on traditional benchmarks (MME) exhibit greater stability.
@@
Setup. We evaluate four different JPEG compression levels: default, 70, 80, and 90. Prompts and data splits are kept fixed. To emphasize rank stability rather than absolute performance, Fig. 6 reports model ranks (1=best) instead of raw accuracies—(a) for BLINK–Relative Depth (RD) and (b) for MME (semantic).
BLINK RD (VPT) vs. MME (semantic). Surprisingly—or perhaps not surprisingly—the rankings in BLINK RD fluctuate considerably, even for closed-source models such as Gemini 2.5 Pro. This sensitivity can be attributed to the fact that VPT tasks demand a more fine-grained understanding of visual tokens than semantic-focused tasks; consequently, even subtle pixel-level changes can influence model predictions. In contrast, Figure 9 shows that MME rankings remain remarkably stable across all four compression settings and nine evaluated models, with only a single ranking inversion compared to the pronounced shifts observed in BLINK RD.
Takeaways. JPEG quality should be standardized—or ideally replaced with lossless formats—for VPT benchmarks; otherwise, compression artifacts alone can meaningfully tilt leaderboard standings when model performances are tightly clustered.
7 Discussions
@@
To partially mitigate these issues, we advocate the following practices:
@@
  • Report uncertainty and rank stability. Accompany accuracies with confidence intervals (e.g., Wilson or bootstrap) and simple rank-stability analyses across markers, splits, or seeds, particularly when the low accuracy tasks have relatively small sample sizes. When intervals overlap substantially, treat models as effectively tied rather than declaring definitive winners.
As a step toward more stable evaluation, we release VPBench, the scaled-up visually prompted benchmark used in our work, together with multiple marker variants and reference evaluation scripts. With performance of common models on VPBench shown in Fig. 3, we show that by reducing variance due to arbitrary design choices, VPBench makes performance differences more reflective of true perceptual ability. While currently limited to depth and correspondence tasks, it provides a foundation for broader analyses and future robustness-aware benchmarks for visual grounding.

### v2

@@ Abstract
A key challenge in evaluating VLMs is testing models’ ability to analyze visual content independently from their textual priors. Recent benchmarks such as BLINK probe visual perception through visual prompting, where questions about visual content are paired with coordinates to which the question refers, with the coordinates explicitly marked in the image itself. While these benchmarks are an important part of VLM evaluation, we find that existing models are surprisingly fragile to seemingly irrelevant details of visual prompting: simply changing a visual marker from red to blue can completely change rankings among models on a leaderboard. By evaluating nine commonly-used open- and closed-source VLMs on two visually prompted tasks, we demonstrate how details in benchmark setup, including visual marker design and dataset size, have a significant influence on model performance and leaderboard rankings. These effects can even be exploited to lift weaker models above stronger ones; for instance, slightly increasing the size of the visual marker results in open-source InternVL3-8B ranking alongside or better than much larger proprietary models like Gemini 2.5 Pro. We further show that low-level inference choices that are often ignored in benchmarking, such as JPEG compression levels in API calls, can also cause model lineup changes. These details have substantially larger impacts on visually prompted benchmarks than on conventional semantic VLM evaluations. To mitigate this instability, we curate existing datasets to create VPBench, a larger visually prompted benchmark with 16 visual marker variants. We open-source VPBench and our analysis framework at: https://lisadunlap.github.io/vpbench/.
Figure 1: Small, seemingly irrelevant changes in visual prompting dramatically alter VLM predictions. Left: Qwen2.5-VL accuracy under different visual marker variants on BLINK relative depth. Changes in marker size, shape, color, and label position lead to significant accuracy swings up to 13%. Right: Leaderboard of VPBench Relative Depth. Such variations can reorder leaderboards, with model rankings shifting even when nothing about the underlying task changes.
@@ 1 Introduction
We explore three such sources of evaluation instability across nine modern VLMs on BLINK [9] as well as two more large scale datasets, VPBench-RD and VPBench-SC, that we curate using DA2K [26], and SPair [16]. First, we demonstrate the effect of sample choice: random resampling of image subsets, matched in its modest size to BLINK and drawn from a fixed pool of visually prompted tasks, leads to substantially reordered model rankings, despite the subsets being statistically indistinguishable in size and difficulty. Second, we examine visual prompt formatting. On four visually prompted datasets, we evaluate our models with 16 different marker styles varying in size, shape, color, and label placement. We find that marker style can cause accuracy swings of up to 21% on the same image-question pairs, often causing ranking reversals among state-of-the-art models. Finally, we show that inference-time implementation details which are imperceptible to humans such as JPEG compression further perturb results in statistically significant ways. Additionally, we find that this is specific to visually prompted tasks, as applying the same intervention to more traditional VLM benchmarks does not significantly change the results. We also demonstrate how this fragility can be exploited to “game” leaderboards. For example, strategically selecting the visual marker to be a square instead of a circle causes a weaker model like InternVL3-8B to rank above stronger models like Gemini 2.5 Pro on the BLINK relative depth estimation task.
These findings suggest that much of the variation among model performance reported on vision-language benchmarks comes not from differentiated intrinsic capabilities of grounding language in vision, but from incidental details of prompting, sampling, and implementation. To address this, we release the expanded visually prompted datasets as VPBench. VPBench is a benchmark covering 16 different visual marker variants, and overall boosting the dataset size from 224 samples in BLINK relative depth and semantic correspondence to 35,088 annotated images across relative depth (VPBench-RD) and semantic correspondence (VPBench-SC). Lastly, we provide suggestions on how to more robustly evaluate visually prompted tasks and guidance on when to trust the leaderboard rankings. We release our proposed VPBench along with our inference code, which supports varying visual markers and image compression settings, as a reference inference for stable evaluation at https://lisadunlap.github.io/vpt-fragile/.
@@ 3 Visually Prompted Tasks and Probing Them
Figure 2: Examples of visually prompted tasks. Visually prompted tasks (VPTs) place visual markers in the image to ask questions such as relative depth and semantic correspondence.
@@ 3.1 What are Visually Prompted Tasks?
A visually prompted task explicitly marks regions of an image and asks about relationships within or between images [9]. We pick two typical visually prompted tasks (Figure 2). In semantic correspondence, two images with marked points are shown and the model is asked which marker corresponds to the same object part or region as the reference marker. In relative depth, two locations in an image are marked with “A” and “B,” and the model is asked which is closer to the camera.
@@ 3.2 Experimental Setup
The BLINK dataset [9] systematizes these visually prompted tasks via curated image–question pairs with region-level markings. It contains relative depth and correspondence tasks with high annotation quality. However, we notice that within BLINK, visual markers are not standardized: some examples use dots, others boxes or arrows. When these seemingly cosmetic differences are altered, we notice non-trivial change in model performance. This motivated our investigation into how VLM benchmarks may be fragile given seemingly irrelevant design choices.
As will be detailed in Section 4, BLINK’s small size, approximately 100 examples per split, makes it difficult to separate true performance variation from sampling noise. To reduce this ambiguity and enable more robust analysis, we curate a new benchmark called VPBench. VPBench is sourced from two larger datasets that follow the same visually prompted task format: DA2K [26] and SPair-71K [16], with the former used for relative depth estimation (VPBench-RD) and the latter for semantic correspondence (VPBench-SC). DA-2K, originally introduced for depth annotation, contains thousands of images with dense geometric labels; we repurpose it into a relative-depth task by converting pixel-level depth into pairwise comparisons following BLINK’s protocol. SPair-71K (SPair), designed for semantic correspondence, provides tens of thousands of image pairs with detailed keypoint matches; by adopting BLINK’s prompting style, we frame SPair as a visually prompted correspondence benchmark that bridges traditional CV evaluation with modern VLM usage. We then apply all interventions for the following sections to all 4 datasets.
@@ 4 Statistically Equivalent Sampling Yields Different Results
Why does sampling matter?
This sensitivity arises because BLINK, like many newly-released multimodal benchmarks, deliberately targets unsaturated capabilities—i.e., tasks where performance levels are far from the ceiling, typically in the 30–60% range rather than the 80–90% common in relatively mature VLM tasks. Consequently, the accuracy variation in visually prompted (VP) tasks is significantly higher than in knowledge-focused benchmarks like MME (see Figure 5). At such accuracy levels and sample sizes, even a few items can noticeably shift results; for instance, in BLINK, a 3% accuracy change corresponds to only three additional correct responses, a margin that could easily arise by chance in a multiple-choice format. This variance is reflected in the confidence intervals in Figure 3. We later highlight this impact by showing how results fluctuate as the dataset size decreases.
@@ 5 Visual Marker Styles Shuffle Leaderboards
We investigate whether the style of the visual marker used in a prompt influences model performance, and whether it can change the relative ordering of models’ accuracy. This question is motivated by an observation of inconsistent marker styles in the BLINK benchmark dataset [9]. The visual prompts BLINK employs as part of the question context occasionally switch in color or text positioning and we notice, for example, that simply switching a marker’s color from red to blue questions lead to measurable changes in a model’s accuracy on BLINK, suggesting current models are over-reliant on specific visual cues. In this section we create a set of marker variants and study their effects on model performance. We report results on VPBench in the main paper and report results for the BLINK subsets in the Appendix.
@@
Figure 6: Small marker changes cause large, model-specific accuracy shifts and rank shuffles. Rank (top) and accuracy (bottom) variations on (a) VPBench-RD relative depth and (b) VPBench-SC semantic correspondence across marker variants (five representative variants shown; full 16 in Appendix). Rows are models and columns are marker styles; each cell reports the change in accuracy relative to the default BLINK marker, which uses a red circle with the label above. Positive values indicate improvements and negative values indicate degradations.
@@
Based on the most common marker seen in BLINK relative depth and semantic correspondence, we define a default marker style to be a small red circular marker with a numeric label placed above. We then create 16 alternative styles (commonly seen in visual prompting literature) spanning 6 categories: color (blue instead of red), shape (square instead of circle), marker size (larger marker, radius increased from the default to size 10), text size (larger/smaller font size), text label (1/2 VS A/B), and text position (label moved to below the marker). Figures 6 and 6 display the change in accuracies and rankings of each model on relative depth and semantic correspondence tasks.
Figure 7: Impact by marker type. Mean absolute accuracy shift on VPBench-RD induced by each marker variant. Changes to font scale and text placement result in the largest accuracy changes.
@@
To determine whether observe accuracy differences between markers are statistically significant rather than due to sampling variability, we draw $B$ bootstrap replicates $\mathcal{D}^{s}$ by sampling $N$ items with replacement. On each replicate, we compute accuracies with the default and an alternative marker and take the paired difference
@@
Figures 6 and Figure 6 show the model accuracy on the default visual marker (red circle, text at the top) along with the delta in performance seen across 16 different visual markers of varying sizes, shapes, colors, and positions. We see that if you change the visual marker, both the absolute accuracy and the relative rankings of models change, which can completely change the overall takeaways. For example, Llama 4 Scout outperforms Gemini 2.5 Flash under the default red marker for relative depth, but under the blue marker condition their ranks reverse, with Gemini 2.5 Flash scoring as the highest performing model. We also see cases of incredibly large differences in accuracy, such as Llama 4 Scout and InternVL for semantic correspondence, where making the text size of the marker smaller results in a 10+% drop in accuracy and dropping Llama from rank 3 to rank 6.
@@
Figure 7 displays the average magnitude of accuracy difference across each marker for the relative depth task, altering text size or label position generally produces larger effects on accuracy, which color and shape of the marker having a less but still noticeable impact. Changing the marker’s color (red to blue) has a large impact on certain models, hinting that those models might be overfit to the color distribution of markers seen in their training or the benchmark (since BLINK’s default is red, some models may have learned to specifically attend to red circles as a prompt cue). Changing the marker’s size and shape or changing the text marker from A/B to 1/2 has a measurable but less prominent effect on accuracy. Notably, there is not a single marker style that was universally best or worst for all models. These idiosyncratic responses point to each model having its own biases in visual prompt processing. Similar plots for semantic correspondence can be found in the Appendix.
@@
We explicitly demonstrate that visually prompted leaderboards can be gamed by jointly selecting a marker style and an i.i.d. split that favor a particular model (Figure 8). On BLINK relative depth, for example, we lower InternVL3 from rank 4 to rank 8 by changing the marker to a square. Conversely, increasing the marker font size raises its rank to 3, placing it above Gemini 2.5 Pro. These results show that, without standardized marker conventions, visually prompted evaluations can give misleading impressions of model ability.
Figure 8: Manipulating leaderboards for InternVL3-8B. Performance comparison: optimizing for InternVL3-8B’s ranking on BLINK relative depth. Leaderboard rank can shift by benign visual prompt choices rather than changes in model capability.
@@ 6 Imperceptible differences matter as well
Figure 9: Change in rank seen across JPEG compression rates. Individual dots with no lines indicate models whose rank remains constant across all compression levels. JPEG compression quality substantially reorders model rankings on visually-dependent tasks (BLINK RD), whereas rankings on traditional benchmarks (MME) exhibit greater stability.
@@
Setup. We evaluate four different JPEG compression levels: default, 70, 80, and 90. Prompts and data splits are kept fixed. To emphasize rank stability rather than absolute performance, Figure 9 reports the change in model ranks for BLINK–Relative Depth and MME. Singular dots mean the model did not change ranks while horizontal lines indicate a change in ranks over different compression rates. The rankings in BLINK RD fluctuate considerably, even for closed-source models such as Gemini 2.5 Pro and GPT-4.1. This sensitivity can likely be attributed to the fact that VPT tasks demand a more fine-grained understanding of visual tokens than semantic-focused tasks; consequently, even subtle pixel-level changes can influence model predictions. In contrast, MME rankings remain relatively stable across all four compression settings and nine evaluated models, with only a single ranking inversion compared to the pronounced shifts observed in BLINK RD. Given this, we recommend that JPEG quality is be standardized or ideally replaced with lossless formats for VPT benchmarks.
7 Discussion
@@
To mitigate these issues, we suggest the following:
@@
  • Report uncertainty and rank stability. Accompany accuracies with confidence intervals and simple rank-stability analyses across markers, splits, or seeds, particularly when the low accuracy tasks have relatively small sample sizes. When intervals overlap substantially, treat models as tied rather than declaring definitive winners.
As a step toward more stable evaluation, we release VPBench, the scaled-up visually prompted benchmark used in our work, together with multiple marker variants and reference evaluation scripts. With performance of common models on VPBench shown in Fig. 3, we show that by reducing variance due to arbitrary design choices, VPBench makes performance differences more reflective of perceptual ability. While currently limited to depth and correspondence tasks, it provides a foundation for broader analyses and future robustness-aware benchmarks for visual grounding.